OpenStudy (anonymous):

What is the Least Common Denominator of the fraction 2 over x plus the fraction x over x plus 1 plus the fraction x plus 1 over 3 x ?

OpenStudy (doc.brown):

I'm thinking, multiply all the denominators together \[\frac{2}{x}+\frac{x}{x+1}+\frac{x+1}{3x}\]\[3x^{3}+3x^{2}?\]

OpenStudy (anonymous):

@doc.brown These are the options: 3x + 1 3(x2 + 1) 3(x + 3) 3x(x + 1)

OpenStudy (doc.brown):

I found a way to multiply x, x+1, and 3x by things to get 3x(x+1) x (3(x+1) = 3x(x+1) (x+1) (3x) = 3x(x+1) 3x (x+1) = 3x(x+1) I can't find what to multiply x+1 with to get 3x+1 x+1 with to get 3(x2 + 1) or x with to get 3(x + 3)